OpenStudy (anonymous):

If f(x)=√9x find f′(x).

OpenStudy (bahrom7893):

f(x) = 3sqrt(x) = 3x^(1/2)

OpenStudy (bahrom7893):

f'(x) = (3/2)x^(-1/2) = 3/(2sqrt(x))

OpenStudy (anonymous):

how did u get f(x) = 3sqrt(x)

OpenStudy (anonymous):

youre allowed to sqrt the 9 even though its conected to x?